# Soliton Amplification in the Korteweg-de Vries Equation by Multiplicative Forcing

### v1

@@ 7 Proof of main result
Assume S1 and the choices (7.1) and (5.1) for S2 and S3. If C4 is satisfied and C5 holds for some $T>0$, then for each $\epsilon,\gamma>0$ the bound
@@
$\left|\partial_{t}\log({\alpha}(t))+\tfrac{2}{3}\epsilon f(\gamma t)\right|+\left|\partial_{t}\Omega(t)-\tfrac{2}{3}\epsilon\frac{f(\gamma t)}{c^{1/2}(t)}\right|\leq C_{6}\big{(}w_{\infty}^{-1/2}\epsilon e^{-\gamma t}\|v(t)\|_{L^{2}_{w_{\infty}}}+\|v(t)\|^{2}_{L^{2}_{w_{\infty}}}\big{)}.$
@@
For each $\epsilon,E>0$ and $\overline{v}_{*}\in H^{2}\cap H^{1}_{w}$ we now introduce the sets
@@
where $C_{8}$ is the constant from Proposition 5.2 and we recall that the weight-function $\mathbf{w}(t)=\big{(}w_{-}(t),w_{+}(t)\big{)}$ is defined in (5.3). We then define $T_{\mathrm{max}}(\epsilon,\gamma,\overline{v}_{*})\in[0,\infty]$ as
@@
Recalling that $E=\epsilon/\gamma$, the key ingredient towards establishing Theorem 1.1 is to show that $T_{\mathrm{max}}(\epsilon,\gamma,\overline{v}_{*})=\infty$.
@@
and consider an initial condition $\overline{v}_{*}\in H^{2}\cap H^{1}_{w}$ that satisfies $\|\overline{v}_{*}\|^{2}_{H^{1}}+\|\overline{v}_{*}\|^{2}_{H^{1}_{w}}\leq\delta_{1}\gamma$. Then the conditions in Theorem 1.1 are met, and we note that
@@
Suppose, for sake of contradiction, that $T_{\mathrm{max}}<\infty$. By construction of $O_{1}$, we have
$\sup_{t\in[0,T_{\mathrm{max}})}e^{\gamma t}\|{v}(t)\|_{H^{1}_{\mathbf{w}(t)}}\leq C_{8}\big{(}\|\overline{v}_{*}\|_{H^{1}_{w}}+\|\overline{v}_{*}\|_{H^{1}}+\epsilon^{1-2p}\big{)}\leq\min\{\delta_{6}w_{\infty}^{1/2},1\}$
for $\delta_{1}$ sufficiently small. Using Lemma 7.1, we obtain
@@
$\leq$ $C_{6}C_{8}w_{\infty}^{-1/2}(\sqrt{2}E_{\mathrm{max}}\delta_{1}+\delta_{1})+4C_{6}C_{8}^{2}\delta_{1}$
@@
$\leq$ $C_{15}\delta_{1}^{2}+C_{15}C_{8}(\sqrt{2}E_{\mathrm{max}}\delta_{1}+\delta_{1})+4C_{15}C_{8}^{2}\delta_{1}.$
@@
$K_{\epsilon,\gamma}(T_{\mathrm{max}}-q)\leq$ $E_{\mathrm{max}}\delta_{1}+\delta_{1}+2C_{8}^{2}(\delta_{1}^{2}+\delta_{1})$
$+\sup_{s\in[0,T_{\mathrm{max}})}\big{(}\|{v}(s)\|_{H^{1}}+e^{\gamma s}\|{v}(s)\|_{H^{1}_{\mathbf{w}(s)}}\big{)}<\delta_{8}$
@@
We may then apply Proposition 5.2 to obtain
@@
Choosing $q<r$ shows that $T_{\mathrm{max}}$ is not maximal, allowing us to conclude that indeed $T_{\mathrm{max}}=\infty$.
@@ Appendix A Decompositions
Let $c>0$ and $0<w<\sqrt{c}$, then $\mathcal{L}_{c}$ is the generator of a $C_{0}$-semigroup on $H^{s}_{w}$ for any real $s$. For all $\beta>0$ which satisfy $\beta<w(c-w^{2})$, there exists a constant $C>0$ such that for all $g\in L_{w}^{2}$, $t>0$ and $k\in\{0,1\}$ we have
@@
Let $c_{*}>0$ and $w\in(0,\sqrt{c_{*}})$, then there exist constants $\delta_{2},C_{2}>0$ such for each $\overline{v}_{*}\in H^{1}_{w}$, the bound $\|\overline{v}_{*}\|_{L^{2}_{w}}\leq\delta_{2}$ implies that there exist unique parameters $c_{0}>0$ and $\xi_{0}\in{\mathbb{R}}$ such that
@@ Appendix B Time-varying weights
we note that for $s\in[0,\delta]$ we may compute
@@
$\frac{{\alpha}(t+s)}{{\alpha}(t)}\leq\frac{(e^{2C_{6}E_{\mathrm{max}}(2+\delta_{6})e^{1/2}})^{-e^{-\gamma(t+s)}}}{(e^{2C_{6}E_{\mathrm{max}}(2+\delta_{6})e^{1/2}})^{-e^{-\gamma(t+s/2)}}}=\frac{w_{-}(t+s)}{w_{-}(t+s/2)}.$
@@
$w_{+}(t)-w_{+}(t+\delta/2)\geq$ $\tfrac{\delta}{2}|w_{+}^{\prime}(t+\delta/2)|$
$=$ $\tfrac{\delta}{2}\gamma\log(\tfrac{w_{\mathrm{min}}}{w})(\tfrac{w_{\infty}}{w})^{1-e^{-\gamma(t+\delta/2)}}e^{-\gamma(t+\delta/2)},$

### v2

@@ 7 Proof of main result
Assume S1 and the choices (7.1) and (5.1) for S2 and S3. If C4 is satisfied and C5 holds for some $T>0$, then for each $\epsilon,\gamma>0$, the bound
@@
$\left|\partial_{t}\log({\alpha}(t))+\tfrac{2}{3}\epsilon f(\gamma t)\right|+\left|\partial_{t}\Omega(t)-\tfrac{2}{3}\epsilon\frac{f(\gamma t)}{c^{1/2}(t)}\right|\leq$ $C_{6}w_{\infty}^{-1/2}\epsilon e^{-\gamma t}\|v(t)\|_{L^{2}_{w_{\infty}}}$
$+C_{6}\|v(t)\|^{2}_{L^{2}_{w_{\infty}}}.$
@@
For each $\epsilon,E>0$ and $\overline{v}_{*}\in H^{2}\cap H^{1}_{w}$, we now introduce the sets
@@
where $C_{8}$ is the constant from Proposition 5.2, and we recall that the weight-function $\mathbf{w}(t)=\big{(}w_{-}(t),w_{+}(t)\big{)}$ is defined in (5.3). We then define $T_{\mathrm{max}}(\epsilon,\gamma,\overline{v}_{*})\in[0,\infty]$ as
@@
Recalling that $E=\epsilon/\gamma$, the key ingredient toward establishing Theorem 1.1 is to show that $T_{\mathrm{max}}(\epsilon,\gamma,\overline{v}_{*})=\infty$.
@@
and consider an initial condition $\overline{v}_{*}\in H^{2}\cap H^{1}_{w}$ that satisfies $\|\overline{v}_{*}\|^{2}_{H^{1}}+\|\overline{v}_{*}\|^{2}_{H^{1}_{w}}\leq\delta_{1}\gamma$. Then, the conditions in Theorem 1.1 are met, and we note that
@@
Suppose, for the sake of contradiction, that $T_{\mathrm{max}}<\infty$. By construction of $O_{1}$, we have
$\sup_{t\in[0,T_{\mathrm{max}})}e^{\gamma t}\|{v}(t)\|_{H^{1}_{\mathbf{w}(t)}}\leq C_{8}\big{(}\|\overline{v}_{*}\|_{H^{1}_{w}}+\|\overline{v}_{*}\|_{H^{1}}+\epsilon^{1-2p}\big{)}\leq C_{8}(\sqrt{2}+1)\delta_{1},$
and
$\sup_{t\in[0,T_{\mathrm{max}})}\gamma^{-1}e^{2\gamma t}\|{v}(t)\|^{2}_{H^{1}_{\mathbf{w}(t)}}\leq 3C_{8}^{2}\gamma^{-1}\big{(}\|\overline{v}_{*}\|_{H^{1}_{w}}^{2}+\|\overline{v}_{*}\|_{H^{1}}^{2}+\epsilon^{2-4p}\big{)}\leq 6C_{8}^{2}\delta_{1}.$
In particular,
$\sup_{t\in[0,T_{\mathrm{max}})}e^{\gamma t}\|{v}(t)\|_{H^{1}_{\mathbf{w}(t)}}\leq\min\{\delta_{6}w_{\infty}^{1/2},1\}$
for $\delta_{1}$ sufficiently small. Thus, we may use Lemma 7.1 to obtain
@@
$\leq$ $C_{6}C_{8}w_{\infty}^{-1/2}(\sqrt{2}E_{\mathrm{max}}\delta_{1}+\delta_{1})+6C_{6}C_{8}^{2}\delta_{1}$
@@
$\leq$ $C_{15}\delta_{1}^{2}+C_{15}C_{8}(\sqrt{2}E_{\mathrm{max}}\delta_{1}+\delta_{1})+6C_{15}C_{8}^{2}\delta_{1}.$
@@
$K_{\epsilon,\gamma}(T_{\mathrm{max}}-q)\leq$ $\sup_{s\in[0,T_{\mathrm{max}})}\big{(}\|{v}(s)\|_{H^{1}}+e^{\gamma s}\|{v}(s)\|_{H^{1}_{\mathbf{w}(s)}}+\gamma^{-1}e^{2\gamma s}\|{v}(s)\|^{2}_{H^{1}_{\mathbf{w}(s)}}\big{)}$
$+E_{\mathrm{max}}\delta_{1}+\delta_{1}<\delta_{8}$
@@
Having established a priori control on $K_{\epsilon,\gamma}$, we may apply Proposition 5.2 to obtain
@@
Choosing $q<r$ shows that $T_{\mathrm{max}}$ is not maximal, allowing us to conclude that, indeed, $T_{\mathrm{max}}=\infty$.
@@ Appendix A Decompositions
Let $c>0$ and $0<w<\sqrt{c}$. Then, $\mathcal{L}_{c}$ is the generator of a $C_{0}$-semigroup on $H^{s}_{w}$ for any real $s$. For all $\beta>0$ which satisfy $\beta<w(c-w^{2})$, there exists a constant $C>0$ such that, for all $g\in L_{w}^{2}$, $t>0$ and $k\in\{0,1\}$, we have
@@
Let $c_{*}>0$ and $w\in(0,\sqrt{c_{*}})$. Then, there exist constants $\delta_{2},C_{2}>0$ such that, for each $\overline{v}_{*}\in H^{1}_{w}$, the bound $\|\overline{v}_{*}\|_{L^{2}_{w}}\leq\delta_{2}$ implies that there exist unique parameters $c_{0}>0$ and $\xi_{0}\in{\mathbb{R}}$ such that
@@ Appendix B Time-varying weights
we note that for $s\in[0,\delta]$, we may compute
@@
$\frac{{\alpha}(t+s)}{{\alpha}(t)}\leq\frac{(e^{2C_{6}E_{\mathrm{max}}(2+\delta_{6})e^{1/2}})^{-e^{-\gamma(t+s)}}}{(e^{2C_{6}E_{\mathrm{max}}(2+\delta_{6})e^{1/2}})^{-e^{-\gamma(t+s/2)}}}=\frac{w_{-}(t+s)}{w_{-}(t+s/2)}.\qed$
@@
$w_{+}(t)-w_{+}(t+\delta/2)\geq\tfrac{\delta}{2}|w_{+}^{\prime}(t+\delta/2)|=\tfrac{\delta}{2}\gamma\log(\tfrac{w_{\mathrm{min}}}{w})(\tfrac{w_{\infty}}{w})^{1-e^{-\gamma(t+\delta/2)}}e^{-\gamma(t+\delta/2)},$